8. A campaign researcher would like to estimate the proportion of registered voters that will vote for his candidate. He runs a poll each week asking registered voters their candidate preference and calculates a 95% confidence interval to estimate the true proportion of registered voters that will vote for his candidate. On week 39 his confidence interval was (0.52, 0.56). On week 40 (the last week of the campaign) he wants to conduct one final poll. He would like his 95% confidence interval to have a margin of error of.01. Use the previous weeks information to determine the minimum sample size needed to achieve the researchers goals for estimation. (A) 96 (B) 9604 (C) 9543 (D) 11187 (E) 9589
Answer #1

here estimate proportion =(0.52+0.56)/2 =0.54

here margin of error E = 0.01
for95% CI crtiical Z          = 1.9600
estimated proportion=p= 0.5400
required sample size n =         p*(1-p)*(σ/E)2= 9543.00

option C is correct